16 g of an ideal gas SOx occupies 5.6L.at STP.The value of x is
 

Grade:10

4 Answers

sree lakshmi
3 years ago
since it is an ideal gas the volume of 1 mol gas is 22.4L
mass of 5.6L gas is 16g(given)
then the mass 22.4 L =(22.4×16)/5.6 
                                  =64g
the mass of SOx is 64g
ie 32+x16=64
x16=32
x=2 
the molecule is SO
[mass of sulpher =32g 
mass of oxygen atom =16g]
adilakshmigogada 27
3 years ago
 
since it is an ideal gas the volume of 1 mol gas is 22.4L
mass of 5.6L gas is 16g(given)
then the mass 22.4 L =(22.4×16)/5.6 
                                  =64g
the mass of SOx is 64g
ie 32+x16=64
x16=32
x=2 
the molecule is SO
[mass of sulpher =32g 
mass of oxygen atom =16g
Parv Joshi
3 years ago
 
ince it is an ideal gas the volume of 1 mol gas is 22.4L
mass of 5.6L gas is 16g(given)
then the mass 22.4 L =(22.4×16)/5.6 
                                  =64g
the mass of SOx is 64g
ie 32+x16=64
x16=32
x=2 
the molecule is SO
[mass of sulpher =32g 
mass of oxygen atom =16g]
irshad hussain
2 years ago

5.6 l is occupied by 16 g of gas

22.4 l is occupied by (16/5.6) × 22.4 g = 64 g

Gram molecular mass = 64 g

Now mass of SOx= 32+ 16x

ATP

32 + 16x= 64

16x = 32

x = 2

The gas is SO2